pH of 10^-8 M solution of HCL in water is between 6 and 7.how 

Grade:11

1 Answers

Atul Nagar
5 years ago
When 1x10^-8 M HCl is added to water, the [H+] increases thus pH decreases.Neutral water contains [H+]= 10^-7M
Total [H+]= (1×10^-7+1×10^-8)
= 1.1×10^-7M
pH = -log[H+]= 7- log1.1 = 7 - 0.0414
pH = 6.9586
 
Askiitian Member
BTech Chem